Lane invested $14,000, part at 15% and part at 2%. If the total inerest at the end of the year is $1,060, how much did she invest at 15%?
Answer by ptaylor(2198)

Let x=amount invested at 15%
Then 14000-x=amount invested at 2%
Interest(I) equals Principal(P) times Rate(R) times Time(T) or I=PRT

Now we know that total interest ($1060) equals interst at 15% (0.15x) plus interest at 2% (0.02)(14000-x). So our equation to solve is:
1060=0.15x+0.02(14000-x) get rid of parens
1060=0.15x+280-0.02x subtract 280 from both sides
1060-280=0.15x+280-280-0.02x collect like terms
780=0.13x divide both sides by 0.13
x=$6000-------------------------------amount invested at 15%
14000-x=14000-6000=$8000--------------amount invested at 2%
CK
6000(0.15)+8000(0.02)=1060
900+160=1060
1060=1060

Hope this helps----ptaylor